OpenStudy (ksaimouli):

4-x^2=ax^2-1 where x=-1

OpenStudy (anonymous):

just plug in -1 so 4-(-1)^2=a(-1)^2-1

OpenStudy (anonymous):

4-1=a-1

OpenStudy (anonymous):

3=a-1

OpenStudy (anonymous):

4=a

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

just replace the 'x' in the equation with the -1. if you need help with it..just go to www.myalgebra.com its a really cool tool for math of all sorts